OpenStudy (anonymous):

if a-b is 6 more than c+d and a+b is 3 less than c-d then a-c is?

OpenStudy (gorv):

a-b=c+d+6 a-c=d+6+b ---------(1) and a+b=c-d-3 a-c=-d-3-b ---------(2) add 1 and 2 2a-2c=3 a-c=3/2

OpenStudy (gorv):

@robertfrost

OpenStudy (anonymous):

a-b-6=c+d (Equation 1) a+b+3=c-d (eqt 2) so now isolate: a-c=b+d+6 (eqt 1) a-c=-b-d-3 (eqt 2) so now sub each one into each other eqt 1= eqt 2 a-c=a-c b+d+6=-b-d-3 2b+2d+9 i'm stuck... this is as far as i got

OpenStudy (gorv):

dont euate just add them up

OpenStudy (gorv):

@dg98